36.
36. (8 pts) Calculate r(t) if a(t) = -32k, v(0) = 31 - 2 + k and r(0) = 5j + 2k.
Answer #1

QIHE -328 dy - - taſ. a due rasket Integrate both sides V= (1-32+ k + Gita Viol: 31 2jt 7 (=3 (25+2 13 = 1 +13) * VHF si-i +